OpenStudy (anonymous):

Given the domain {-5, 2, 3}, what is the range for the relation 16x + 4y = 12? A) {-1, 5, 6} B) {-23, 5, 9} C) {-17, 11, 15} D) {23, -5, -9} thanks!:)

OpenStudy (dumbcow):

solve for y, then plug in x-values from domain

OpenStudy (anonymous):

Answer =option(D) {23 , -5 , -9}

OpenStudy (anonymous):

16x + 4y = 12 solve for y 16x + 4y = 12 -16x -16x ------------------- 4y = 12 - 16x (divide by 4 on both side) y = 3 - 4x Now plug the x-values(that is domain value) to find y (that is range) plug x=-5 plug x=2 plug x=3 y=3-4(-5) y=3-4(2) y=3-4(3) y=3+20 y=3-8 y=3-12 y=23 y=-5 y=-9 Therefore,Range={23,-5,-9}

OpenStudy (anonymous):

I hope that you got it